An electron is confined in a cubical box of each side 1 Å. Calculate the energies of the electron in

the ground state and the first excited state


1
Expert's answer
2021-05-06T07:40:21-0400

Answer

Energy for ground state

"E=\frac{h^2}{8mL^2}\\E=\frac{(6.63\times10^{-34})^2}{8(9.1\times10^{-31})(10^{-10})^2}\\
=37.737eV"



"E_2=4E=150.95eV"


"E_3=9E=339.639eV"